a). An insulated container is used to hold 40.4 g of water at 29.6 °C. A sample of copper weighing 15.6 g is placed in a dry test tube and heated for 30 minutes in a boiling water bath at 100.0°C. The heated test tube is carefully removed from the water bath with laboratory tongs and inclined so that the copper slides into the water in the insulated container. Given that the specific heat of solid copper is 0.385 J/(g·°C), calculate the maximum temperature of the water in the insulated container after the copper metal is added. °C?


b) Calculate the amount of energy needed to convert 10.1 g of water at 29.0 °C to steam at 100.0°C.


1
Expert's answer
2021-02-03T07:15:22-0500

Solution:


a)

Let T be the equilibrium temperature (in °C) to be found.


(4.184 "\frac{J}{g}" ·°C) x (45.3 g) x (T - 19.0) °C = (189.535 T - 3601.17) J gained by the water


(0.385 "\frac{J}{g}" ·°C) x (11.7 g) x (100.0 - T) °C = (450.45 - 4.5045 T) J lost by the copper


Set the two expressions for the heat gained/lost equal to each other:


189.535 T - 3601.17 = 450.45 - 4.5045 T

Solve for T algebraically:

T ≈ 20.9 °C


b)

"Q=mc\Delta T=(10.1g)(4.184\;\frac{J}{g*C})(100-29)C=3000.34\;J=3kJ"